OpenStudy (anonymous):

What is the quotient written in scientific notation? 3x10^20 over 4x10^5

OpenStudy (anonymous):

300,000,000,000,000,000,000,/400,000

OpenStudy (trojanpoem):

(3/4) * 10^(20-5) When you divide you simply subtract the powers.

OpenStudy (anonymous):

0.75x10^25

OpenStudy (anonymous):

i mean 15

OpenStudy (anonymous):

0.75x10^15

OpenStudy (trojanpoem):

3/4 = 3/2/2 = 15/2/10 = 7.5/10 = 0.75 * 10^15